# Towards Sharp Minimax Risk Bounds for Operator Learning

### v1

@@ 7 Conclusion
We developed a minimax theory for learning uniformly bounded Lipschitz operators from noisy input-output data when both the inputs and outputs may be infinite dimensional. Our general upper and lower bounds cover fixed and random designs and accommodate both Hilbert-valued Gaussian noise and Gaussian white noise. Our results make precise how statistical difficulty is governed by the eigenvalues of the covariance of $\mu$. Our results also recover classical finite-dimensional Lipschitz rates as a special case.
There are several open problems that remain. Most importantly, our lower bounds are not sharp for algebraically decaying eigenvalues and, more generally, for exponential decay with $0<\beta<1$. A key technical obstacle is the $d$-dependence in the current lower-bound argument. Removing it would likely lead to a sharp characterization in these regimes. In the algebraic case, we conjecture that the true minimax rate is polylogarithmic in the sample size, with an exponent determined by the decay rate of the eigenvalues. Even in regimes where our bounds match, such as exponential and double-exponential eigenvalue decay, identifying the optimal constant in the exponent appears nontrivial. Finally, for very rapidly decaying eigenvalues (such as double-exponential decay) one can achieve ‘nearly’ algebraic performance. Beyond Lipschitz operators, extending the framework to higher-regularity operator classes (e.g., the full scale of Hölder and Besov classes) is a natural direction for future work.

### v2

@@ 7 Conclusion
We developed a minimax theory for operator learning from noisy input-output data when both the inputs and outputs may be infinite dimensional. Our general upper and lower bounds cover fixed and random designs and accommodate both Hilbert-valued Gaussian noise and Gaussian white noise. Our results make precise how statistical difficulty is governed by the eigenvalues of the covariance of $\mu$. Our results also recover classical finite-dimensional rates as a special case.
There are several open problems that remain. In the case of Lipschitz operators, most importantly, our lower bounds are not sharp for algebraically decaying eigenvalues and, more generally, for exponential decay with $0<\omega<1$. A key technical obstacle is the $d$-dependence in the current lower-bound argument. Removing it would likely lead to a sharp characterization in these regimes. In the algebraic case, we conjecture that the true minimax rate is polylogarithmic in the sample size, with an exponent determined by the decay rate of the eigenvalues. Even in regimes where our bounds match, such as exponential and double-exponential eigenvalue decay, identifying the optimal constant in the exponent appears nontrivial. Finally, for very rapidly decaying eigenvalues (such as double-exponential decay) one can achieve ‘nearly’ algebraic performance. Beyond Lipschitz operators, we showed that higher-order (Hölder) regularity does not change the bounds. This shows that imposing higher, but still finite, regularity cannot overcome the curse of sample complexity in operator learning. Extending this framework to other operator classes (e.g., holomorphic, where lower bounds are not known, or Besov classes, where neither upper nor lower bounds are known) is another natural direction for future work.
Acknowledgements
BA acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) through grant RGPIN/2470-2021. GM acknowledges support from the Hausdorff Center for Mathematics (HCM) in Bonn, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813.